# Brain Foundation Models: A Survey on Advancements in Neural Signal Processing and Brain Discovery

### v1

@@
Xinliang Zhou,  Chenyu Liu,  Zhisheng Chen,  Kun Wang,  Yi Ding,  Ziyu Jia,  and Qingsong Wen Equally Contribution Corresponding AuthorsManuscript received March 1, 2025.
Brain foundation models (BFMs) have emerged as a transformative paradigm in computational neuroscience, offering a revolutionary framework for processing diverse neural signals across different brain-related tasks. These models leverage large-scale pre-training techniques, allowing them to generalize effectively across multiple scenarios, tasks, and modalities, thus overcoming the traditional limitations faced by conventional artificial intelligence (AI) approaches in understanding complex brain data. By tapping into the power of pretrained models, BFMs provide a means to process neural data in a more unified manner, enabling advanced analysis and discovery in the field of neuroscience. In this survey, we define BFMs for the first time, providing a clear and concise framework for constructing and utilizing these models in various applications. We also examine the key principles and methodologies for developing these models, shedding light on how they transform the landscape of neural signal processing. This survey presents a comprehensive review of the latest advancements in BFMs, covering the most recent methodological innovations, novel views of application areas, and challenges in the field. Notably, we highlight the future directions and key challenges that need to be addressed to fully realize the potential of BFMs. These challenges include improving the quality of brain data, optimizing model architecture for better generalization, increasing training efficiency, and enhancing the interpretability and robustness of BFMs in real-world applications. Ethical, regulatory, and standardization considerations also remain key aspects to be addressed as these models are applied in clinical and healthcare settings. By bridging the gap between neuroscience and AI, BFMs have the transformative potential to revolutionize brain disease diagnosis, neurorehabilitation strategies, and cognitive enhancement techniques. Their ability to integrate complex neural signals and make sense of vast amounts of data can provide solutions for more effective and individualized healthcare. This work aims to serve as a foundational reference for researchers, practitioners, and policymakers, highlighting the transformative impact of BFMs on both our understanding of the human brain and the broader implications they hold for the future of brain-related technologies.
@@ I Introduction
Neural signal processing has advanced dramatically over the past century, driven by progress in neuroscience and computational methods, as shown in Fig. 1. Early research focused on brain recordings to study sleep, epilepsy, and cognition, leading to the adoption of statistical techniques like principal component analysis (PCA) [1] and independent component analysis (ICA) [2] for reducing dimensionality and isolating meaningful signal components amidst noisy environments. By identifying core features hidden within linearly-structured neural data, PCA and ICA facilitated significant breakthroughs in both basic science and clinical applications, for example, identifying distinct electroencephalogram (EEG) frequency bands and distinguishing epileptic spike patterns. Traditional machine learning models, such as support vector machines (SVMs), random forests, and k-nearest neighbors (KNN) [3], further advanced these developments by providing more robust classification and regression frameworks, addressing challenges ranging from automated seizure detection to more applications, such as distinguishing emotional states from EEG. However, traditional machine learning methods were limited by their reliance on handcrafted features, sensitivity to variations in data quality across recording setups, and limited ability to model the nonlinear and dynamic nature of brain activity. As experimental paradigms became more diverse, spanning motor imagery, visual evoked potentials, and cognitive load tasks, the shortcomings of manual feature engineering and the poor adaptability of conventional machine learning methods became increasingly evident.
@@
These challenges sparked increasing interest in foundation models (FMs), large-scale systems pretrained on broad data distributions to learn universal representations. The remarkable success of FMs in natural language processing (NLP) and computer vision (CV), exemplified by models like BERT [7], GPT [8], and Segment Anything (SAM) [9], demonstrated the power of leveraging large, diverse datasets and self-supervised objectives to capture complex statistical patterns. By pretraining on text, images, or video from various sources, these models learned generalized features that could be adapted to a wide array of downstream tasks, often outperforming traditional models in similar applications. Inspired by this, researchers began exploring the potential of FMs for neural signals, hypothesizing that their ability to capture complex dependencies could benefit EEG and functional magnetic resonance imaging (fMRI) analysis. With pretraining techniques like masked signal modeling, FMs showed promise for significant performance improvements in tasks like mental state classification, motor imagery, and neurodegenerative disease diagnosis. However, directly applying these FMs to neural data presents fundamental challenges. Compared to text or images, neural signals exhibit greater spatiotemporal complexity and often have a lower signal-to-noise ratio (SNR). Additionally, recordings can vary significantly across individuals and are often subject to strict ethical constraints, including patient privacy and institutional review protocols. Moreover, effective brain research and clinical translation require simultaneous support for decoding, diagnosis, and dynamic network modeling, an integrated approach that extends beyond the focus of most existing FMs.
@@
Figure 1: Progression of neural signal processing techniques from traditional methods to BFMs. Especially, BFMs significantly enhance the precision and scope of large-scale neural data analysis, promoting breakthrough discoveries in neuroscience. In 2022, the BFM paradigm has been continuously followed by research scholars, leading to a boom in new models such as BrainLM, Brant, LaBraM, and NeuroLM in recent years. This highlights rapid advancements and increasing diversity in brain research methodologies.
@@
To bridge this gap, Brain Foundation Models (BFMs) offer a paradigm that unifies neural signal processing through large-scale, neurocentric pretraining and cross-scenario, cross-task capabilities. Unlike general-purpose FMs, which typically rely on massive text or image corpora, BFMs aim to incorporate large-scale datasets of neural signals, such as EEG or fMRI recordings spanning thousands of subjects and numerous hours. The BFMs approach is underpinned by three guiding principles: (1) pretraining tailored to the dynamics of neural data, (2) enabling zero- or few-shot generalization across distinct scenarios, tasks, and modalities, reflecting the broad variability of brain states and experimental conditions in neuroscience, and (3) embedding ethical AI mechanisms that address the sensitive nature of human neural recordings, such as federated learning frameworks to protect data privacy and rigorous anonymization protocols to safeguard participant identities.
@@
  • Finally, we provide a forward-looking roadmap, highlighting open research challenges, including data integration and quality, novel training strategies, multimodal architectural design, interpretability, and ethical and regulatory considerations that must be addressed to fully realize the potential of BFMs.
@@ II-A1 Definition
BFMs refer to foundational models built using deep learning and neural network technologies pretraining on large-scale neural data designed to decode or simulate brain activity. These models aim to capture and understand the complex patterns in neural signals, thereby advancing neuroscience exploration and enabling brain disease diagnosis and treatment. BFMs integrate multimodal brain signal processing (e.g., EEG and fMRI), biological principles, and artificial intelligence techniques to extract deep neural activity patterns from large-scale data and multidimensional features. This allows for a multi-angle and precise interpretation of brain function.
BFMs can be classified into three categories:
  • Pretrained Only: These models are constructed by pretraining on large-scale brain signal datasets to develop a robust feature extraction capability, which can be applied to general brain activity analysis and disease screening. After pretraining, these models are ready for deployment without the need for further fine-tuning.
  • Pretrained with Fine-tuning: These models are pretrained on extensive brain signal data and then fine-tuned for specific applications. This approach is used for tasks such as brain disease diagnosis, cognitive state assessment, and other specialized applications.
  • Pretrained with Interpretability for Brain Discovery: These models combine pretraining with interpretability techniques, such as perturbation analysis, to simulate and explore key biological mechanisms in brain activity. The goal of this approach is not only to provide high-precision diagnostic support but also to drive brain discovery by analyzing the neural underpinnings of brain functions.
@@ II-A2 Framework
The BFM framework, as shown in Fig. 2, provides a clear view of how BFMs and their applications are constructed. The framework is structured around three key components: large-scale neural data, pretraining (with or without fine-tuning), and interpretability analysis. These components work synergistically to enable BFMs to decode and simulate brain activity across different scenarios and tasks.
  • Large-Scale Neural Data: At the core of the BFM framework is the utilization of large-scale datasets of neural signals, including EEG, fMRI, and other brain activity recordings. These datasets are crucial as they provide the necessary diversity and volume to allow the model to learn generalized representations of brain function. The data often span across thousands of subjects and hours of brain activity, helping the model identify universal patterns that are applicable to a broad range of brain states and conditions. A larger, more varied dataset improves the BFM’s ability to generalize across experimental conditions, which is crucial for real-world applications.
@@
    The BFM framework allows for two paths in model training: pretraining with or without fine-tuning.
@@ II-C1 Data Processing
Remark: Despite initial progress, the exploration of data processing strategies in current BFMs remains in its early stages, with substantial room for improvement. While existing methods have demonstrated effectiveness in standardizing neural data for large-scale modeling, they continue to face fundamental challenges in handling variations in temporal structures and brain region configurations, limiting their adaptability across different experimental paradigms and individuals.
A key limitation of current BFMs is the fixed time window strategy, which struggles to accommodate the diverse temporal structures of different experimental paradigms. Predefined window lengths often fail to capture the dynamics of neural activity: in event-related potential analysis, excessively long windows introduce noise, diluting critical patterns, while in sustained attention tasks, overly short windows truncate essential information, disrupting neural continuity. Beyond task-specific concerns, inter-individual variability further complicates this issue, as subjects exhibit differing neural response latencies and processing speeds. The mismatch between static window lengths and the inherent variability of neural signals introduces biases, ultimately limiting the adaptability and robustness of current BFMs.
Another major challenge is handling inconsistent brain region counts across datasets. Existing approaches either take the intersection of recorded regions, as in CBRAMOD [16], sacrificing information for compatibility, or take the union, as in EEGPT [17], increasing computational complexity. These rigid standardization strategies fail to account for the natural variability in brain recordings, limiting their ability to capture functional connectivity patterns and region-specific dynamics. Furthermore, treating brain regions as discrete channels prevents biologically meaningful learning, as current models lack explicit constraints to encode structural relationships. This underscores the need for more flexible representations that can dynamically adapt to varying brain region configurations while preserving essential physiological information.
@@ II-C2 Training Strategy
It is a fundamental component of BFMs, as it defines how models extract meaningful representations from complex neural signals. Given the challenges posed by brain data, such as high variability, individual differences, and task-dependent dynamics, effective training paradigms are crucial for enabling BFMs to generalize across diverse experimental conditions. Existing BFMs primarily adopt two broad training approaches: reconstruction-based learning and contrastive learning.
@@
Methodologies: The first training strategy for BFMs is reconstruction-based self-supervised learning (SSL). Its core principle is to mask or corrupt a portion of the input signal and require the model to recover the missing part, thereby learning the intrinsic structure of the data. This approach is widely applied in brain signal modeling as it effectively extracts both local and global features from temporal sequences, enhancing the model’s understanding of neural activity. Reconstruction-based methods can generally be categorized into two types:
  • Masked Prediction: A portion of the signal is randomly masked, and the model is trained to reconstruct the missing information based only on the visible parts.
  • Autoregression: The model is constrained to use only historical data to predict future neural signals in a sequential manner.
Masked prediction aims to learn global representations from incomplete input data. From a neuroscience perspective, this approach aligns with the redundant coding principle of the brain, wherein neural activity in certain regions can often be inferred from surrounding information. For instance, EEG signals exhibit strong temporal correlations, allowing masked time segments or channels to be reconstructed from the remaining data. Moreover, the brain employs redundant representation mechanisms, in which multiple brain regions collaboratively encode the same cognitive or motor states. This property enables masked learning strategies to capture latent functional patterns in neural activity. For example, BrainBERT [18] adopts a time-frequency masking strategy, where randomly selected segments of the EEG spectrograms are obtained by the short-time Fourier transform or the Superlet transform and subsequently reconstructs them. It employs a Transformer-based architecture, leveraging self-attention mechanisms to extract contextual information from the unmasked regions and predict the missing spectral features. Brain-JEPA further refines this masking strategy by introducing spatiotemporal joint masking. Unlike BrainBERT, which randomly masks specific time-frequency points, Brain-JEPA employs cross-region, cross-time, and dual masking strategies on fMRI data, ensuring that the model generalizes effectively under various masking conditions. Additionally, instead of directly reconstructing the raw signal, Brain-JEPA predicts features in a latent representation space, a design that mitigates noise interference and improves the stability of learned representations.
@@
In contrast, autoregressive methods predict future neural signals solely based on past observations, aligning with the brain’s temporal integration mechanism, where current neural states largely depend on preceding activity. For example, the transmission of signals across brain regions often follows inherent time delays and specific neural functions exhibit consistency across multiple temporal scales. Thus, autoregressive models can effectively capture the dynamic evolution of brain activity and learn long-range dependencies in neural signals. Unlike masked prediction, autoregressive methods emphasize modeling long-term dependencies, which is crucial for understanding causal relationships and information flow within the brain. BrainLM [19], for instance, employs a Transformer-based autoregressive framework for fMRI data, where the model sequentially predicts future time steps based on past neural activity. This training strategy enables the model to effectively learn long-range temporal dependencies, improving its ability to represent brain dynamics. NeuroLM [20] further advances this approach by introducing a multi-channel autoregressive mechanism, which integrates information across EEG channels to enhance joint modeling. Moreover, it incorporates a causal autoregressive framework, enforcing directional constraints in the prediction process, thereby improving the model’s capability to capture EEG temporal patterns with greater fidelity.
BFMs adopt another training strategy based on contrastive learning, which aims to construct positive and negative sample pairs to learn more discriminative representations by pulling similar samples closer while pushing dissimilar ones apart. Contrastive learning methods can be generally categorized into two types:
  • Intra-Brain Modality Contrastive Learning: Contrastive learning within brain data, including comparisons within the same modality or across different brain modalities, to integrate multi-source neural signals and construct more comprehensive representations of neural activity.
  • Brain-Heterogeneous Modality Contrastive Learning: Contrastive learning between brain data and heterogeneous modalities (such as text or images) to represent neural information in alternative modalities and enhance its interpretability.

### v2

@@
Xinliang Zhou,  Chenyu Liu,  Zhisheng Chen,  Kun Wang,  Yi Ding,  Ziyu Jia,  and Qingsong Wen Equally Contribution Corresponding AuthorsManuscript received March 1, 2025, revised June 1, 2025, and accepted on July 15, 2025.
Brain foundation models (BFMs) represent a transformative paradigm in computational neuroscience that leverages large-scale pre-training on diverse neural signals to achieve robust generalization across tasks, modalities, and experimental contexts. This survey establishes the first comprehensive definition and framework for BFMs, systematically examining their construction, core methodologies, and applications. We present key approaches for data processing and training strategies alongside diverse applications spanning brain decoding and scientific discovery. Through critical analysis of recent methodological innovations, we identify fundamental challenges that must be addressed to realize the full potential of BFMs, including advancing data quality and standardization, optimizing model architectures, improving training efficiency, and enhancing interpretability. By bridging the gap between neuroscience and artificial intelligence, BFMs present unprecedented opportunities to revolutionize brain research, clinical diagnostics, and therapeutic interventions. This survey serves as a foundational reference for researchers and practitioners seeking to understand and advance this emerging field.
@@ I Introduction
Neural signal processing has advanced dramatically over the past century, driven by progress in neuroscience and computational methods, as shown in Fig. 1. Early research focused on brain recordings to study sleep, epilepsy, and cognition, leading to the adoption of statistical techniques like principal component analysis (PCA) [1] and independent component analysis (ICA) [2] for reducing dimensionality and isolating meaningful signal components amidst noisy environments. By identifying core features hidden within linearly-structured neural data, statistical methods facilitated significant breakthroughs in both basic science and clinical applications, for example, identifying distinct electroencephalogram (EEG) frequency bands and distinguishing epileptic spike patterns. Traditional machine learning models, such as support vector machines (SVMs), random forests, and k-nearest neighbors (KNN) [3], further advanced these developments by providing more robust classification and regression frameworks. They addressed challenges ranging from automated seizure detection to more applications, such as distinguishing emotional states from EEG. However, traditional machine learning methods were limited by their reliance on handcrafted features, sensitivity to variations in data quality across recording setups, and limited ability to model the nonlinear and dynamic nature of brain activity. As experimental paradigms became more diverse, spanning motor imagery, visual evoked potentials, and cognitive load tasks, the shortcomings of manual feature engineering and the poor adaptability of conventional machine learning methods became increasingly evident.
@@
These challenges sparked increasing interest in foundation models (FMs)[7], large-scale systems pretrained on broad data distributions to learn universal representations. The remarkable success of FMs in natural language processing (NLP) and computer vision (CV), exemplified by models like BERT [8], GPT [9], and Segment Anything (SAM) [10], demonstrated the power of leveraging large, diverse datasets and self-supervised objectives to capture complex statistical patterns. By pretraining on text, images, or video from various sources, these models learned generalized features that could be adapted to a wide array of downstream tasks, often outperforming traditional models in similar applications. Inspired by this, researchers began exploring the potential of FMs for neural signals, hypothesizing that their ability to capture complex dependencies could benefit EEG and functional magnetic resonance imaging (fMRI) analysis. With pretraining techniques like masked signal modeling, FMs showed promise for significant performance improvements in tasks like mental state classification, motor imagery, and neurodegenerative disease diagnosis. However, directly applying these FMs to neural data presents fundamental challenges. Compared to text or images, neural signals exhibit greater spatiotemporal complexity and often have a lower signal-to-noise ratio (SNR). Additionally, recordings can vary significantly across individuals and are often subject to strict constraints, including patient privacy and institutional review protocols. Moreover, effective brain research and clinical translation require simultaneous support for decoding, diagnosis, and dynamic network modeling, an integrated approach that extends beyond the focus of most existing FMs.
@@
Figure 1: Evolution of neural signal processing from traditional methods to BFMs. Panels (a) to (c) illustrate the progression from statistical techniques (e.g., PCA, ICA) and machine learning (e.g., SVM, KNN) to deep learning (e.g., CNN, LSTM) and BFMs. BFMs enhance the precision and scope of large-scale neural data analysis, enabling breakthroughs in neuroscience. Since 2022, BFMs like BrainLM and LaBraM have driven rapid advancements and methodological diversity in brain research.
@@
To bridge this gap, Brain Foundation Models (BFMs) offer a paradigm that unifies neural signal processing through large-scale, neurocentric pretraining and cross-scenario, cross-task capabilities. Unlike general-purpose FMs, which typically rely on massive text or image corpora, BFMs aim to incorporate large-scale datasets of neural signals, such as EEG or fMRI recordings spanning thousands of subjects and numerous hours. The BFMs approach is underpinned by three guiding principles: (1) pretraining tailored to the dynamics of neural data, (2) enabling zero- or few-shot generalization across distinct scenarios, tasks, and modalities, reflecting the broad variability of brain states and experimental conditions in neuroscience, and (3) embedding AI mechanisms that address the sensitive nature of human neural recordings, such as federated learning frameworks to protect data privacy and rigorous anonymization protocols to safeguard participant identities.
@@
  • Finally, we provide a forward-looking roadmap, highlighting open research challenges, including data integration and quality, training strategies, multimodal architectural design, interpretability and regulatory considerations that must be addressed to fully realize the potential of BFMs.
@@ II-A1 Definition
BFMs refer to foundation models built using deep learning and neural network technologies pretraining on large-scale neural data designed to decode or simulate brain activity. These models aim to capture and understand the complex patterns in neural signals, thereby advancing neuroscience exploration and enabling brain disease diagnosis and treatment. BFMs integrate multimodal brain signal processing (e.g., EEG and fMRI), biological principles, and artificial intelligence techniques to extract deep neural activity patterns from large-scale data and multidimensional features. This allows for a multi-angle and precise interpretation of brain function.
BFMs can be classified into three categories based on their training and application strategies. Pretrained-only models are constructed by pretraining on large-scale brain signal datasets to develop a robust feature extraction capability, which can be applied to general brain activity analysis and disease screening. After pretraining, these models are ready for deployment without the need for further fine-tuning. Pretrained with fine-tuning models are pretrained on extensive brain signal data and then fine-tuned for specific applications. This approach is used for tasks such as brain disease diagnosis, cognitive state assessment, and other specialized applications. Pretrained with interpretability for brain discovery models combine pretraining with interpretability techniques, such as perturbation analysis, to simulate and explore key biological mechanisms in brain activity. This goal is not only to provide high-precision diagnostic support but also to drive brain discovery by analyzing the neural underpinnings of brain functions.
@@ II-A2 Framework
The BFMs framework, as shown in Fig. 2, provides a clear view of how BFMs and their applications are constructed. The framework is structured around three key components: large-scale neural data, pretraining (with or without fine-tuning), and interpretability analysis. These components work synergistically to enable BFMs to decode and simulate brain activity across different scenarios and tasks.
  • Large-Scale Neural Data: At the core of the BFMs framework is the utilization of large-scale datasets of neural signals, including EEG, fMRI, and other brain activity recordings. These datasets are crucial as they provide the necessary diversity and volume to allow the model to learn generalized representations of brain function. The data often span across thousands of subjects and hours of brain activity, helping the model identify universal patterns that are applicable to a broad range of brain states and conditions. A larger, more varied dataset improves the BFM’s ability to generalize across experimental conditions, which is crucial for real-world applications.
@@
    The BFMs framework allows for two paths in model training: pretraining with or without fine-tuning.
@@ II-C1 Data Processing
Remark: Unlike foundational models in other domains, brain data exhibits differences in position encoding across various dimensions. In the temporal dimension, neural signals exhibit relatively stable temporal patterns, making fixed-position encoding more suitable to enhance generalization and avoid overfitting. In the spatial dimension, functional connectivity and topological structures between brain regions vary significantly across individuals and task contexts, making learnable position encoding more appropriate to accommodate diverse spatial relationships and task-related patterns. Building on this, the rotation-based position encoding adopted by current mainstream foundational models offers a solution that balances stability and expressive power. It retains the advantages of fixed encoding while introducing relative position modeling capabilities, making it suitable for modeling long-term dependencies in neural signals.
Summary of Data Processing: Data processing for BFMs is still in its early stages and faces challenges that limit its adaptability. The fixed time window strategy often does not match the varied timing of neural activity or individual differences, which can introduce biases and reduce model robustness. Managing inconsistent brain region configurations across datasets is another major issue; methods like regional intersections (e.g., CBRAMOD [17]) can cause information loss, while unions (e.g., EEGPT [18]) can increase the computational load. These inflexible standardization techniques, along with treating brain regions as separate channels without considering their structural relationships, also hinder biologically meaningful learning. Therefore, there is a strong need for more flexible and adaptive representations that can manage variations in temporal structure and brain region configurations while keeping essential physiological information.
@@ II-C2 Training Strategy
It is a fundamental component of BFMs, as it determines how models extract meaningful representations from complex and noisy neural signals. Given the high variability, individual differences, and task-dependent dynamics inherent in brain data, effective training paradigms are essential to enable BFMs to generalize across diverse experimental conditions.
Most BFMs follow a two-stage training paradigm: pretraining and fine-tuning. During pretraining, models typically leverage self-supervised learning (SSL) to learn generalized neural signal representations from large-scale and diverse datasets without requiring labeled data. Some BFMs are then fine-tuned on specific downstream tasks using supervised learning, while others are designed for direct application post-pretraining (pretrained-only models, such as NeuroLM[19] ).
Methodologies: SSL is central to the pretraining phase and operates by defining proxy tasks that enable models to learn useful features from unlabeled data. In BFMs, SSL objectives primarily fall into two categories: reconstruction-based learning, where models learn to recover missing parts of the input from the available context, and contrastive learning, where models learn to distinguish between similar and dissimilar pairs of brain signals.
Formally, let $x_{i}\in\mathbb{R}^{d}$ represent the input brain signal, where $d$ denotes the dimensionality of the data (for example, the number of EEG channels). The goal is to learn a function $f$ such that the learned representation $f\left(x_{i}\right)$ captures the underlying patterns in the signal. A common SSL objective is to minimize the reconstruction error or to maximize the similarity between different parts of the input data, which is expressed mathematically as:
$\mathcal{L}_{\text{SSL }}=\alpha\sum_{i}\mathcal{L}_{\text{reconstruction }}\left(x_{i},\hat{x}_{i}\right)+\beta\mathcal{L}_{\text{contrastive }}\left(f\left(x_{i}\right),f\left(x_{j}\right)\right)$ (2)
where $\hat{x}_{i}$ represents the model’s prediction for $x_{i}$, and $\mathcal{L}_{\text{reconstruction }}$ is a reconstruction loss (e.g., Mean Squared Error), which ensures that the model learns to predict missing parts of the signal. The contrastive loss $\mathcal{L}_{\text{contrastive }}$ is used to make the model learn similar representations for similar inputs and dissimilar representations for different ones. The coefficients $\alpha\in\{0,1\}$ and $\beta\in\{0,1\}$ act as indicators to select the desired loss components, where $(\alpha,\beta)\neq(0,0)$.
  • If $\alpha=1$ and $\beta=0$, the SSL objective focuses solely on reconstruction.
  • If $\alpha=0$ and $\beta=1$, the SSL objective focuses solely on contrastive learning.
  • If $\alpha=1$ and $\beta=1$, both reconstruction and contrastive learning objectives are jointly optimized.
@@
Reconstruction-Based Learning: Its core principle is to mask or corrupt a portion of the input signal and require the model to recover the missing part, thereby learning the intrinsic structure of the data. This approach is widely applied in brain signal modeling as it effectively extracts both local and global features from temporal sequences, enhancing the model’s understanding of neural activity. Reconstruction-based methods can generally be divided into two main types: masked prediction and autoregression. Masked prediction involves randomly masking segments of the neural signal, training the model to reconstruct the missing information based solely on the unmasked portions. Autoregression constrains the model to rely exclusively on historical data, enabling it to predict future neural signals in a sequential manner.
Masked prediction aims to learn global representations from incomplete input data. From a neuroscience perspective, this approach aligns with the redundant coding principle of the brain, wherein neural activity in certain regions can often be inferred from surrounding information. For instance, EEG signals exhibit strong temporal correlations, allowing masked time segments or channels to be reconstructed from the remaining data. If $X$ represents the input brain signal segments and $\tilde{X}$ is the version with certain segments masked (denoted by a set of masked indices $M$), and $f_{\theta}$ is the BFM with parameters $\theta$ that attempts to reconstruct the masked segments $\hat{X}_{masked}$ from the unmasked segments $X_{unmasked}$, the objective is often to minimize a reconstruction loss. For example, using Mean Squared Error (MSE) for continuous signals, this can be formulated as:
$L_{\text{recon}}=\frac{1}{|M|}\sum_{i\in M}(X_{i}-\hat{X}_{i})^{2}$ (3)
where $X_{i}$ is the original value of a masked element and $\hat{X}_{i}$ is its predicted value by the model $f_{\theta}(X_{\text{unmasked}})$. The specific nature of $X$ (e.g., raw signal values, spectral features) and the masking strategy can vary.
In this context, masking can be performed in various ways, depending on the task at hand:
  • Temporal Masking: Specific time windows in the neural signal (e.g., EEG or fMRI) are masked. The model learns to predict the missing segments from surrounding temporal information. This is particularly useful in modeling dynamic brain activity over time.
  • Spatial Masking: In fMRI or EEG data, certain spatial regions (such as brain regions or channels) are masked, and the model is tasked with reconstructing or predicting the activity in these regions based on the remaining parts of the signal.
For example, BrainBERT [20] adopts a time-frequency masking strategy, where randomly selected segments of the EEG spectrograms are obtained by the short-time Fourier transform or the Superlet transform and subsequently reconstructs them. It employs a Transformer-based architecture, leveraging self-attention mechanisms to extract contextual information from the unmasked regions and predict the missing spectral features. Brain-JEPA further refines this masking strategy by introducing spatiotemporal joint masking.
In contrast, autoregressive methods predict future neural signals solely based on past observations, aligning with the brain’s temporal integration mechanism, where current neural states largely depend on preceding activity. If a neural signal sequence is denoted as $X=(x_{1},x_{2},...,x_{T})$, an autoregressive BFM $f_{\theta}$ aims to predict the signal $x_{t}$ given the past $k$ observations $(x_{t-k},...,x_{t-1})$. The training objective typically involves minimizing the prediction error across the sequence, often defined as:
$L_{\text{autoreg}}=\sum_{t=k+1}^{T}\mathcal{D}(x_{t},f_{\theta}(x_{t-k},...,x_{t-1}))$ (4)
@@
where $\mathcal{D}$ represents a suitable divergence or distance function (e.g., MSE for continuous signals or cross-entropy if signals are discretized into tokens). BrainLM [21], for instance, employs a Transformer-based autoregressive framework for fMRI data, where the model sequentially predicts future time steps based on past neural activity. NeuroLM further advances this approach by introducing a multi-channel autoregressive mechanism, which integrates information across EEG channels to enhance joint modeling. Moreover, it incorporates a causal autoregressive framework, enforcing directional constraints in the prediction process, thereby improving the model’s capability to capture EEG temporal patterns with greater fidelity.
Remark: Different modalities of brain data are suitable for different SSL strategies. Masked prediction is more suitable for modalities such as fMRI. These modalities have high spatial resolution, and brain regions often exhibit significant spatial co-activation patterns under the same cognitive state. By masking the signals of certain brain regions and requiring the model to perform reconstruction, the model can be guided to fully utilize the spatial dependencies in the data. Autoregression is more suitable for modalities with strong temporal continuity, such as EEG. These signals have high temporal resolution, reflecting the rapid dynamic changes in brain activity and exhibiting distinct temporal evolutionary characteristics. By predicting future signals based on historical time segments, the model can effectively capture the causal structure and temporal dependencies of neural activity.
Contrastive Learning: It aims to construct positive and negative sample pairs to learn more discriminative representations by pulling similar samples closer while pushing dissimilar ones apart. This approach has been highly successful in visual and multimodal domains, notably exemplified by models like CLIP [22], which aligns image and text representations. Many contrastive learning methods aim to maximize the agreement between different ’views’ of the same underlying brain state or activity. Given an anchor sample $x_{i}$ (e.g., a segment of a brain signal), a positive sample $x_{i}^{+}$ (e.g., an augmented version of $x_{i}$ or a concurrently recorded related signal from another modality), and a set of $N-1$ negative samples $x_{k}$ (dissimilar signals), a common objective is the InfoNCE loss (or a variant):
$L_{\text{contrastive}}=-\mathbb{E}\left[\log\frac{\exp(\text{sim}(f_{\theta}(x_{i}),f_{\theta}(x_{i}^{+}))/\tau)}{\exp(\text{sim}(f_{\theta}(x_{i}),f_{\theta}(x_{i}^{+}))/\tau)+\sum_{k=1}^{N-1}\exp(\text{sim}(f_{\theta}(x_{i}),f_{\theta}(x_{k}))/\tau)}\right]$ (5)
where $f_{\theta}$ is the BFM encoder that maps inputs to a representation space, $\text{sim}(\cdot,\cdot)$ denotes a similarity function (e.g., cosine similarity) between representations, and $\tau$ is a temperature hyperparameter. This loss encourages the model to learn representations that are invariant to certain augmentations (in intra-modality learning) or that align across different modalities (in brain-heterogeneous or inter-brain modality learning). Contrastive learning methods can generally be categorized into two types: intra-brain modality contrastive learning and brain-heterogeneous modality contrastive learning. The former focuses on contrastive objectives within brain data, either within the same modality or across different brain modalities. This approach facilitates the integration of multi-source neural signals and helps build more comprehensive representations of neural activity. The latter, on the other hand, involves contrasting brain data with data from heterogeneous modalities such as text or images. This strategy enables the representation of neural information in alternative modalities, thereby enhancing interpretability and broadening potential applications.